What is the amount of money I will have after a year if I put $5,000 in my Savings Bond account and the bank pays me 3.5% each year?
Answer by stanbon(75887)

What is the amount of money I will have after a year if I put $5,000 in my Savings Bond account and the bank pays me 3.5% each year?
-------
Simple Interest::
A(t) = P(1 + rt)
----
A(1) = 5000(1+0.035*1) = 5000(1.035) = $5175
----------------
Cheers,
Stan H.
-----------------